OpenStudy (lxelle):

solve 8 cos ^4 x-5=0

OpenStudy (yolo_king):

In order to do this, you must first simplify by reordering the terms.

OpenStudy (yolo_king):

\[-5+8\cos^4x=0\]

OpenStudy (yolo_king):

Then we move all the terms containing c to the left, all the other terms to the right.

OpenStudy (yolo_king):

\[-5+5+8\cos^4x=0+5\]

OpenStudy (yolo_king):

Then we combine like terms.

OpenStudy (yolo_king):

-5+5=0

OpenStudy (yolo_king):

\[0+8\cos^4x=0+5\]

OpenStudy (yolo_king):

8cos^4x=0+5

OpenStudy (yolo_king):

Then we divide each side by\[8os^4x\]

OpenStudy (yolo_king):

We would get \[c=0.625o^{-1}s^{-4}x^{-1}\]

OpenStudy (yolo_king):

Normally we would simplify but it is already in simplest form...

OpenStudy (lxelle):

???

OpenStudy (mathmate):

We'll restart from here, up to which is good: 8cos^4x=0+5 so 8cos^4x=5 Divide both sides by 8 \(cos^4(x)=5/8\) take 4th root: \(\rm cos(x)=\pm \dfrac{5}{8}^{1/4}\) The right hand side is numeric and can be evaluated using a calculator (appr.0.9), and the corresponding x=acos(0.9)=0.48 Now examine the cosine function  Assuming x=0.48 (radians) is a solution (first intersection of line & curve), then so are \(\pi-x\), \(\pi+x\), and \(2\pi-x\). These 4 solutions are between 0 and 2pi. They repeat at every interval of 2pi, since cos(x) is periodic. So the general solution is x=\(k\pi \pm 0.48\) (approximate because 0.9 is an approximate answer) where k\(\in\)Z (all integers)